OpenStudy (anonymous):

Solve –2x2 +3x – 9 = 0

OpenStudy (btaylor):

do you know the quadratic formula?

OpenStudy (anonymous):

yah, and i think i have the answer just not sure.... heres my answer: ill draw it!;)

OpenStudy (anonymous):



OpenStudy (anonymous):

is thaat right?

OpenStudy (btaylor):

@Ashley1997 it is a negative x^2. So you have to use the quadratic formula

OpenStudy (campbell_st):

a = -2, b = 3 and c = -9\[x = \frac{-3\pm \sqrt{3^2 - 4 \times(-2)\times(-9)}}{2\times(-2)}\] gives \[x = \frac{3 \pm \sqrt{-63}}{4}=\frac{3\pm 3i \sqrt{7}}{4}\]

OpenStudy (campbell_st):

your answer is correct cassie